OpenStudy (anonymous):

Solve the following equation by transforming it into a perfect square trinomial: x^2 – 2x = 3

OpenStudy (nurali):

x² - 2x = 3 x² - x = 3 + (- 1)² x² - x = 3 + 1 (x - 1)² = 4 x - 1 = 4 x = 4 + 1, x = 5 x = - 4 + 1, x = - 3 Answer: (- 3, 5)

OpenStudy (anonymous):

thanks again :)

OpenStudy (nurali):

Anytime.

OpenStudy (jhannybean):

even though you have to solve it using a square BInomial, you can solve it using the quadratic formula too. \[y=ax^2+bx+c\] formula to follow \[x^2-2x-3=0\]\[(x-3)(x+1)=0\]solve for x \[(x-3)=0\]\[(x-2)=0\]\[x=3,2\]

OpenStudy (jhannybean):

@Nurali "x-1=4" JUST equals x=5. With \[\sqrt{(x-1)^2} = |x-1| , |x-1| \neq x-1\]